### v1

# Second-Generation Planet Formation in Post-AGB Discs: Testing the Role of Gravitational Instability

Abstract
Post-Asymptotic Giant Branch (post-AGB) binary stars are evolved systems that host circumbinary discs formed through mass loss during late-stage binary interaction. Their structural, morphological, kinematic, and chemical similarities to planet-forming discs have motivated the idea that these systems exhibit the characteristics of second-generation protoplanetary discs and may act as sites of “second-generation” planet formation. In this study, we assess whether the disc instability mechanism – a proposed pathway for giant planet formation in some protoplanetary discs—can operate in post-AGB discs. We test gravitational instability as a potential planet formation pathway, motivated by the short lifetimes ($10^{4-5}$years) of post-AGB discs, which would require any planet formation mechanism to operate efficiently on rapid timescales. Using the Toomre criterion under well-motivated assumptions for disc structure and size, mass, and thermal properties, we assess the conditions for gravitational instability. We first benchmark our analytical framework using well-studied protoplanetary disc systems (including HL Tauri, Elias 2-27, GQ Lupi) before applying the same analysis to observed post-AGB discs. We find that post-AGB discs are generally gravitationally stable at present, due primarily to their low masses. The disks were stable against collapse, even when accounting for potentially higher disk mass in the past. In contrast, several protoplanetary discs analysed in the same way show that they likely experienced gravitationally unstable phases early on. Finally, we revisit the planet formation scenario proposed for the post-common envelope system NN Serpentis first carried out by Schleicher and Dreizler and we show that gravitational instability could be feasible under specific, high disc mass assumptions, however we note that their proposed disc mass is unlikely. Overall, our results provide the first systematic theoretical assessment of gravitational instability in post-AGB discs, demonstrating that this mechanism is unlikely to dominate second-generation planet formation in these systems, and underscoring the need to explore alternative pathways—such as core accretion—in future studies.
keywords:
stars: AGB and post-AGB, planets and satellites: formation, protoplanetary disks, instabilities
[…]
1 Introduction
Understanding how and where planets form across different stages of stellar evolution remains one of the central questions in modern astrophysics. The Kepler and Gaia space telescopes have provided precise planet and stellar parameters for thousands of exoplanet systems 2010Borucki; 2018Thompson; 2016Gaia; 2020berger and high angular resolution observations with the Atacama Large Millimeter/ sub-millimeter Array (ALMA) have revealed direct signatures of ongoing planet formation in YSO discs, such as gaps, rings, cavities, and spirals - as seen in systems like PDS 70 Keppler_2019, HL Tauri Booth_2020, AB Aurigae Boccaletti_2020, Elias 2-27 Perez_2016, and HD 169142 Fedele_2017.
Planets have also been detected or suspected around evolved stars. Notably, the first confirmed exoplanets were found orbiting the pulsar PSR B1257+12 — a system now known to host at least three planets 1992Wolszczan. The formation and/or survival pathway leading to planets existing in these extreme environments remains uncertain. For instance, 1993Podsiadlowski proposed that the disc from which pulsar planets formed originated through channels distinct from those forming YSO discs. Another example is the L2 Puppis system, an Asymptotic Giant Branch (AGB) star with an edge-on circumbinary disc and a candidate planetary companion – offering a glimpse into a potential future for our own solar system Kervella_2016. Similarly, NN Serpentis, a post-common envelope binary, exhibits eclipse timing variations interpreted as evidence for two circumbinary planets Beuermann_2010. These planets have been proposed to form from gas falling back following a common envelope phase 2014schleicher, as their low masses suggest that they may not have survived direct interaction with the RGB/AGB progenitor. However, alternative explanations — including magnetic activity cycles (the Applegate mechanism) and other dynamical effects — remain under debate Navarrete_2018; _zd_nmez_2023. Systems such as these point to the intriguing possibility of planet formation in previously unanticipated post-main sequence environments.
Of specific interest to us are the discs around post-asymptotic giant branch (post-AGB) binaries. Post-AGB binaries are systems with a post-AGB star in the late stages of stellar evolution. The spectral energy distributions (SEDs) of the majority of post-AGB binary stars show a near-IR excess indicative of a circumbinary disc 2006deRuyter; 2014Kamath; Kamath_2015; Kluska_2022. The formation mechanisms of these discs are still poorly understood, but are generally presumed to have formed from late stage binary interactions, with more than 80 such discs now catalogued (kamath_2019; Kluska_2022; Andrych_2023, e.g.). The observational properties of these discs have been extensively studied and reveal striking similarities to those of protoplanetary discs around YSOs. For instance, interferometric maps with ALMA and observations of 12CO and 13CO lines Bujarrabal_2015; Gallardo_Cava_2023 have shown that many of these discs exhibit stable (quasi)-Keplerian dynamical properties 2006deRuyter; Gallardo_Cava_2021, with inner cavities in the dust component of the disk which is comparable to the dust sublimation radius in many of these systems Kluska_2019; Corporaal_2023. Furthermore, there is evidence of grain growth and dust grain evolution comparable to that seen in protoplanetary discs 2005deruyter; Scicluna_2020; Kluska_2022; Andrych_2023. Moreover, detailed chemical abundance studies of post-AGB binaries systems reveal a photospheric chemical depletion of refractory elements Kluska_2018, which has been interpreted as possibly being caused by a planet forming a dust-trapping gap in these discs. Photospheric depletions are stronger in post-AGB discs categorised as ”transition discs” mohorian2025tracingchemicaldepletionevolved, supporting the possibility that these discs harbour planets. Finally, the morphology of several ”transition” post-AGB discs includes features such as cavities (see Corporaal2023), with inner radii of up to 7.5 times the sublimation radius. These similarities have led to the hypothesis that post-AGB discs may also support planet formation.
[…]
Planets around main sequence stars are presumed to form primarily via core accretion or gravitational instability (GI) drazkowska2023planetformationtheoryera. Core accretion is a bottom-up process involving the growth of small, grain-sized solids through coagulation and sticking during collisions and other mechanisms, eventually forming planetesimals and planetary embryos that can accrete additional material through gravitational interactions. The alternative theory, gravitational instability, is a top-down process in which massive regions of the disc undergo gravitational collapse, forming bound clumps that can contract and accrete material until planet-sized bodies emerge. This mechanism —similar in nature to the process of star formation— has been explored in seminal works such as 1997boss; 2002mayer. Gravitational instability is typically invoked to explain planet formation in the outer regions of discs, where cooling timescales and disc mass conditions may favour instability and fragmentation.
Gravitational instability can lead to planet formation on timescales comparable to the orbital period of the protoplanetary clump, making it a promising candidate for forming planets within short-lived post-AGB discs with lives of up to $10^{5}$ years Bujarrabal_2016. However, this mechanism also has limitations; most notably, it requires a sufficiently massive disc with the mass concentrated in a compact region for fragmentation to occur. Observationally, most post-AGB discs fall below this threshold. Nevertheless, 2014schleicher applied this mechanism to the NN Ser system, arguing that a relatively massive and compact disc – left behind after the CE phase – could have undergone gravitational fragmentation. Their models produced planet masses comparable to those observed in NN Ser, supporting the plausibility of second-generation planet formation via disc instability in select cases.
In this study, we carry out a thorough analytical analysis to ascertain the possibility of planet formation in post-AGB discs through gravitational instabilities, while leaving core accretion to be investigated in a future study. We have anchored our analysis in a range of YSO systems for which planets, or planet-indicative disc features are reported.
This paper is structured as follows. In Section 2, we introduce the analytical framework used to assess the possibility of planet formation via gravitational instability. In Section 3, we apply this framework to a sample of young stellar object (YSO) discs that are known or suspected to host planets or exhibit signatures of gravitational instability, as an initial test of the theory. Section 4 presents the core focus of this work - an investigation of whether typical post-AGB discs can be gravitationally stable. In Section 5, we explore how the masses and sizes of present-day discs may have evolved over time and assess their early-life stability. Section 6 revisits the analysis by 2014schleicher of the NN Serpentis system, where we attempt to reproduce their calculations using more accurate assumptions about the disk’s temperature distributions.
2 Analytical Framework to Investigate Gravitational Instabilities
In this section, we outline the theoretical framework we have used to evaluate the conditions under which a disc may become gravitationally unstable and fragment. We describe the Toomre stability criterion, cooling efficiency conditions, and various models for the disc surface density and thermal structure. These provide the basis for evaluating whether a given disc structure is susceptible to planet formation by fragmentation.
Gravitational instabilities occur in discs if the self-gravity of the disc is able to overcome the effects of thermal pressure and rotation. This is quantified by the Toomre parameter $Q$ 1964Toomre:
$Q(r)=\frac{c_{\rm s}(r)\Omega(r)}{\pi G\Sigma(r)}\leq 1$ (1)
where $c_{\rm s}$ is the sound speed, $\Omega$ is the Keplerian angular frequency of the disc, $\Sigma$ is the disc’s surface density, $G$ is the gravitational constant , and $r$ is the radial distance from the center of the disc. While the classical Toomre criterion requires $Q\leq 1$ for instability against axisymmetric perturbations, simulations indicate that non-axisymmetric instabilities and spiral structures can emerge for $Q\leq 1.4$–1.7 Helled_2014.
The sound speed is related to the temperature, $T(r)$, by the following expression:
$c_{\rm s}(r)=\sqrt{\eta\frac{k_{\rm B}T(r)}{\mu m_{\rm p}}},$ (2)
where $k_{\rm B}$ and $m_{\rm p}$ are the Boltzmann constant and proton mass, respectively, and we have assumed the adiabatic index $\eta=1$  [^1] We denote the adiabatic index by $\eta$ to avoid confusion with the common use of $\gamma$ as a viscosity exponent in disc models. We adopt $\mu=2$ corresponding to molecular hydrogen gas.
For the optically thick regions of the disc:
$T(r)=\left(\frac{\alpha}{4}\right)^{1/4}\left(\frac{R_{\rm star}}{r}\right)^{1/2}T_{\rm star}$ (3)
Chiang_1997, where $\alpha\approx 0.4R_{\rm star}/r$; $\alpha<<1$ is the grazing angle of the flux of stellar radiation incident on the disc, $R_{\rm star}$ is the photospheric radius of the star and $T_{\rm star}$ is the surface temperature of the star. This gives $T\propto r^{-0.75}$. This temperature profile is valid for optically thick discs in radiative equilibrium.
Using this, we propose two temperature profiles for two regions inside and outside of the sublimation radius defined as:
$R_{\rm subl}=\frac{1}{2}R_{\rm star}\left(\frac{T_{\rm star}}{T_{\rm subl}}\right)^{2},$ (4)
where we assume $T_{\rm subl}=1200$ K is the dust sublimation temperature in post-AGB discs Kluska_2019. Assuming radiative equilibrium, the temperature of the disc in the midplane will be
$T(r)=T_{\rm star}\left(\frac{R_{\rm star}}{2r}\right)^{0.5}\,\text{for}\,r<R_{\rm subl},$ (5)
$T(r)=\frac{A}{r^{0.75}}\,\text{for}\,r>R_{\rm subl}.$
where $A$ is a constant that can be calculated by making use of the fact that the two equations should predict the same temperature, $T_{\rm subl}$, at $r=R_{\rm subl}$:
$A=T_{\rm subl}R_{\rm subl}^{0.75}.$ (6)
Using this we have:
$T(r)=T_{\rm star}\left(\frac{R_{\rm star}}{2r}\right)^{0.5}\,\text{for}\,r<R_{\rm subl},$ (7)
$T(r)=T_{\rm subl}\left(\frac{R_{\rm subl}}{r}\right)^{0.75}\,\text{for}\,r>R_{\rm subl}.$
2.1 Cooling Efficiency
In order for gravitational instabilities to lead to disc fragmentation, the fragments must cool sufficiently rapidly such that thermal pressure does not halt the collapse. This criterion is expressed by the following inequality 2001gammie:
$t_{\rm cool}<\frac{\beta}{\Omega},$ (8)
where $t_{\rm cool}$ is the cooling time and $\beta\approx 3$ is a constant. It has been shown in simulations by 2002mayer that protoplanetary discs with minimum Toomre parameters of $Q_{\rm min}\sim 1.4$ start forming spiral structures which can fragment and form several planets within $\sim 1000$ years, if the cooling is efficient enough. The ratio of the thermal energy of the disc to its emissivity can be used to estimate cooling times Rafikov_2005; Kratter_2010:
$t_{\rm cool}=\frac{c_{s}^{2}\Sigma}{(\eta-1)}\frac{f(\tau)}{\sigma T^{4}},$ (9)
where $\eta=1.4$ for molecular Hydrogen-dominated gas, $\sigma$ is the Stefan–Boltzmann constant, and $f(\tau)=\tau+1/\tau$ is a function to quantify the efficiency of disc cooling, which takes the vertical optical depth to be $\tau=\kappa\Sigma/2$, where $\kappa$ is the opacity.
2.2 The Surface Density Profile
The surface density profile, $\Sigma(r)$, plays a key role in determining the gravitational stability of the disc. A commonly adopted empirical form is a truncated power-law 2014schleicher:
$\Sigma(r)=\Sigma_{0}\left(\frac{r_{\rm out}}{r}\right)^{n},$ (10)
where $r$ is the orbital distance, $r_{\rm out}$ is the outer disc’s radius, $n$ is the power law index that determines the steepness of the surface density profile and $\Sigma_{0}$ is a normalisation constant set by the disc’s mass, $M_{\rm disc}$, $n$, and the outer radius:
$\Sigma_{0}=\frac{(2-n)M_{\rm disc}}{2\pi r_{\rm out}^{2}},$ (11)
which is only valid for $n<2$. This type of function is used to fit the surface densities of both YSO discs 2023Speedie, and post-AGB discs Kluska_2022. For a fixed disc mass, decreasing $r_{\rm out}$ increases the overall surface density, effectively making the disc more compact (Figure 1).
In Section 2.2.2, we introduce an alternative surface density formulation that addresses some of the limitations of this parametrisation, including the assumption of a sharply defined outer edge.
2.2.1 Discs with Inner Cavities
Post-AGB discs have inner cavities not considered by Equations 10. Moreover, Equation 11 gives negative surface density values for power law indices $n\geq 2$ . We renormalise the surface density for discs with a finite inner radius, $r_{\rm in}$, resulting in the following:
$\Sigma_{0}$ $=\frac{(2-n)M_{\rm disc}}{2\pi(r_{\rm out}^{2}-r_{\rm in}^{2}(\frac{r_{\rm out}}{r_{\rm in}})^{n})},\>{\rm for}\>n\neq 2,$ (12)
$\Sigma_{0}$ $=\frac{M_{\rm disc}}{2\pi\ln({r_{\rm out}/r_{\rm in}})r_{\rm out}^{2}},\>{\rm for}\>n=2.$
which can be used for all values of $n$. For $n<2$ and $r_{\rm in}=0$ we recover Equation 11. Physically motivated values for $r_{\rm in}$ include the dust sublimation radius or $\sim$3 times the binary separation, corresponding to the innermost stable orbit in circumbinary systems Artymowicz1994.
Figure 1: Power-law surface density functions with $n=1$ as a function of orbital distance for $M_{\rm disc}=0.15M_{\odot}$ for different values of $r_{\rm out}$, to demonstrate why the surface density is highly sensitive to $r_{\rm out}$.
2.2.2 Surface Density Profiles of Viscous Discs
While power-law models are commonly used to describe disc surface density profiles due to their simplicity, they are not physically motivated and rely on poorly defined parameters such as the outer radius, $r_{\rm out}$. Observationally, different diagnostics yield inconsistent estimates of $r_{\rm out}$ — for instance, optically thin dust continuum observations often indicate smaller disc sizes than optically thick molecular line tracers 2009Andrews.
A more physically motivated model, widely adopted for YSO discs, is based on viscous disc evolution theory. This approach originates from the seminal works of 1974lyndenbell and 1981pringle, and was formulated in a practical surface density profile by 1998hartmann. In this model, the disc’s kinematic viscosity is assumed to follow a radial power law with index $\gamma$: $\nu\propto r^{\gamma}$. Under this assumption, the surface density profile takes the form:
$\Sigma(r)=\Sigma_{0}\left(\frac{r}{r_{\rm ch}}\right)^{-\gamma}\exp\left[-\left(\frac{r}{r_{\rm ch}}\right)^{2-\gamma}\right],$ (13)
where $r_{\rm ch}$ is the characteristic radius, and $\Sigma_{0}$ is the normalisation constant. Unlike the outer radius in power-law models, $r_{\rm ch}$ is well-defined within the viscous evolution framework and avoids the sensitivity associated with defining a sharp truncation.
Following Booth_2020, the normalization constant $\Sigma_{0}$ for this profile can be expressed in terms of the total disc mass $M_{\rm disc}$ and $r_{\rm ch}$ as:
$\Sigma_{0}=(2-\gamma)\frac{M_{\rm disc}}{2\pi r_{\rm ch}^{2}}\exp\left[\left(\frac{r_{\rm in}}{r_{\rm ch}}\right)^{2-\gamma}\right],$ (14)
where $r_{\rm in}$ is the inner radius of the disc, if the disc has one.
This exponentially tapered surface density profile has been successfully applied to fit the data from the observation of many YSO discs; however, it has not yet been widely applied to characterise post-AGB discs.
3 Gravitational Instability in YSO Discs
Before applying our gravitational instability framework to the less explored environment of post-AGB discs, we first test its applicability using a set of well-studied YSO systems. Many of these systems show direct or indirect evidence of planet formation, and some exhibit structures–such as spirals, clumps, or large gaps–that have been associated with gravitational instabilities Cadman2021. By applying our analytical model to these systems, we aim to establish a benchmark for interpreting results in post-main-sequence discs.
The YSO systems we consider are:
  1. 1.
    AB Auriga: a system with a gravitationally unstable disc and a detected planet 2022currie.
  2. 2.
    HL Tauri: a disc showing concentric rings with gaps, possibly indicating embedded planets 2015ALMA; Booth_2020.
  3. 3.
    Elias 2-27: a system with spiral structures and density waves, suggesting possible gravitational instability Paneque_Carre_o_2021.
  4. 4.
    L1448 IRS3B: a triple protostar system embedded in a massive, gravitationally unstable disc with prominent spiral arms and a massive clump, as revealed by ALMA observations Reynolds_2021.
  5. 5.
    GQ Lupi: a T Tauri star with a $10-40M_{\rm Jup}$ companion at $\sim 110AU$. While the current disc mass is relatively low ($\sim$3–4 $M_{\rm Jup}$), the wide separation of the companion suggests it may have formed via gravitational instability during an earlier, more massive phase of the disc Wu_2017. We note that the companion may lie in the brown dwarf regime, though its wide separation and possible disc origin make it relevant to GI-driven formation scenarios.
For comparison, we also include two YSO systems with active planet formation, but without prominent indicators of disc instability:
  1. 1.
    PDS 70: a low-mass disc system with two directly imaged planets Keppler_2019.
  2. 2.
    HD 169142: a system with a disc that shows gaps hinting at two planets, one of which has recently been confirmed through direct imaging and ALMA molecular line observations Fedele_2017; Hammond_2023.
In Table 1 we summarise the relevant properties of these systems used in our analysis. For AB Auriga and L1448 IRS3B systems, we assume a power-law surface density profile (for $\Sigma$, see Section 2.2), following 2023Speedie and Reynolds_2021. For HL Tauri and Elias 2-27, where stellar temperatures and radii are not available, we adopt plausible values consistent with their estimated stellar masses. We did not find an estimation for the inner radius of the disk of PDS-70 in the literature, hence we set its inner radius to $r_{\rm in}=0$.
In Figure 2, we plot the Toomre parameter $Q$ (see Section 2.1) as a function of disc radius for each system. While our focus here is on the Toomre criterion as a first-order diagnostic, we note that gravitational fragmentation also requires rapid cooling (see Section 2.1), which we do not assess in this section. We find that only two of the systems are Toomre-unstable at present (AB Aurigae with the disc mass reported in 2023Speedie, and L1448 IRS3B) indicating that gravitational instability is not generally active under current disc conditions for most of the systems, although some of them such as HL Tauri, GQ Lup, and Elias 2-27 are close to the unstable regime. We note that Toomre analyses have been carried out in the past for HL Tauri Booth_2020 and Elias 2-27 Perez_2016, and the resulting Toomre profiles in these studies match the profiles we found for them. This suggests that planet formation via gravitational instabilities may have occurred in these YSOs earlier in the disc’s evolution. We revisit this possibility in Section 5, where we explore whether these systems could have been gravitationally unstable at earlier stages.
Figure 2: Toomre parameter $Q$ versus disc radius for the YSOs tabulated in Table 1. Only AB Aurigae with the disc mass reported in 2023Speedie, and L1448 IRS3B fall below $Q=1$.
Table 1: Stellar and disc parameters for the benchmark YSO systems used in our gravitational instability analysis (see A for sources).
[…]
Planets detected 1 none [^2] No planets confirmed via direct imaging or RV, though disc structures suggestive of embedded planets exist 2015ALMA none 2 1 none [^3] The mass of the clump observed in L1448 IRS3B is high enough to be a protostar. 1
disc mass, $M_{\rm disc}$ ($\rm M_{\odot}$) $0.0032$, 0.70.7444Two values exist in the literature; see A; in this analysis, we have plotted the Toomre curve of this disc using both disc mass values in Figure 2. $0.2$ $0.08$ $0.003$ $0.008$ $0.29$ $0.0733$
Inner radius, $r_{\rm in}$ (au) $80$ $8.78$ 55 0  [^5] No value available - we set the inner radius to 0. $13$ $0.1$ $1.7$
Characteristic/Outer radius [^6] For all systems except AB Aurigae and L1448 IRS3B, the tabulated value in this row is the characteristic radius. For AB Aurigae and L1448 IRS3B, the tabulated values are their outer radii., $r_{\rm ch}$$r_{\rm in}$ (au) $500$ $80$ $200$ $40$ $100$ $299$ $19.5$
Star mass, $M_{\rm star}$ ($\rm M_{\odot}$) $2.4$ $1.7$ $0.46$ $0.85$ $1.65$ $1.2$ $1.05$
Star temperature, $T_{\rm star}$ (K) $9770$ $4000$ $3850$ $3972$ $7500$ 50005000777Indirectly calculated from the luminosity of the protostar, assuming a blackbody spectrum. $4300$
Star radius, $R_{\rm star}$ ($\rm R_{\odot}$) $2.5$ $6.9$ $2.3$ $1.26$ $1.6$ $2.5$ $1.7$
$\gamma$ 0.2-0.2 11 11 11 0.21-0.21
$n$ 11 $1.2$
4 Gravitational Instability in Typical Post-AGB Discs
Having used our GI framework on well-studied YSO systems, we now apply the same methodology to the post-AGB binary population.
Estimates of post-AGB disc gas masses, based on ALMA observations of molecular line emission (e.g. CO), typically fall in the range of $10^{-4}$$10^{-2}$ M (Bujarrabal_2013; Gallardo_Cava_2021, e.g.). In contrast, disc dust masses inferred from radiative transfer modelling of near- and mid-infrared interferometric data and spectral energy distributions (SEDs) can be significantly higher—reaching up to 0.10.10.20.2 M in a few individual cases (Kluska_2018; Corporaal_2023, e.g.). These estimates are sensitive to assumptions about grain properties, vertical structure, and disc inclination, and are often subject to model degeneracies. Interpreting such high dust masses with a canonical gas-to-dust ratio of $\sim$100  [^8] A key uncertainty in interpreting both gas and dust measurements is the gas-to-dust ratio in evolved systems. While a ratio of $\sim$100 is commonly assumed in protoplanetary discs, post-AGB discs may have undergone significant chemical evolution or gas depletion. The true ratio remains unconstrained, and thus dust-based masses may overestimate, and gas-based masses may underestimate, the total disc mass. would imply total disc masses of several solar masses, in tension with dynamical constraints and gas-based mass estimates Hillen_2015; Kluska_2018. To account for the uncertainties in total disc mass estimates—arising from the discrepancy between gas and dust measurements and the unknown gas-to-dust ratio—we adopt a conservative range of 0.010.010.10.1 M for the total (gas + dust) disc mass. We assess gravitational stability across this range using a power-law surface density profile, consistent with standard characterisations of post-AGB discs from observations (Section 2.2). Typical truncation radii lie between 100 and 500 au; for our fiducial case, we adopt $r_{\rm out}=100$ au to represent the more compact end of the distribution, where surface densities are highest and instability most likely.
We assume a central binary mass of $M_{1}+M_{2}=1.5$ M, consistent with typical post-AGB binaries (Oomen_2018, e.g.), and adopt $R_{\rm star}=100$ R and $T_{\rm star}=5,000$ K for the post-AGB star. Assuming a binary separation of $1$ au, we set the disc inner radius to the expected dynamical truncation radius of $r_{\rm in}=3$ au following Artymowicz1994.
In Figure 3, we show the Toomre parameter $Q$ as a function of radius for discs with varying surface density power-law indices. Post-AGB discs appear generally stable across this range. For power-law indices $n\geq 2$, the surface density declines steeply with radius, and stability increases accordingly due to the rapidly decreasing mass at large radii. However, at smaller radii, discs with steeper profiles ($n>2$) can approach marginal instability.
Although power-law indices with $n>2$ appear in some observational fits to post-AGB discs Kluska_2022, such profiles are not physically plausible for a self-gravitating disc without an inner cavity, as the enclosed mass would diverge. To reflect the observed structure and ensure physical consistency, we adopt a power-law (Equation 10) to parametrise the surface densities of post-AGB disks following how they are typically modelled in literature, and we also include an inner cavity in our analysis (Section 2.2, Equation 12) to avoid this divergence and to remain consistent with the typical observed structure of post-AGB discs. This is in contrast with how we have parametrised YSO disks (Equation 13) in the preceding section.
Even under compact, high-density configurations, the resulting $Q$ values remain above the fragmentation threshold in most regions, indicating that typical post-AGB discs are gravitationally stable under present-day conditions.
Figure 3: The Toomre parameter $Q$ as a function of radius for typical post-AGB discs, assuming $M_{1}+M_{2}=1.5$ M, $R_{\rm star}=100$ R, $T_{\rm star}=5,000$ K, $r_{\rm in}=3$ au, and $r_{\rm out}=100$ au as a lower limit of post-AGB disc truncation radii. The index $n$ is the power-law index of the surface density profile for each of the curves. Even in this relatively compact scenario, we see that the discs are generally stable.
4.1 Testing the Stability of the Massive Disc in IRAS 08544-4431
IRAS 08544-4431 is a post-AGB system with an infrared excess, indicating the presence of an optically thick circumbinary disc. Its inner rim lies at the dust sublimation radius Corporaal_2023, consistent with a disc in an early evolutionary stage.  [^9] The inner rim of the dust disc observed in the infrared may not coincide with the inner radius of the gas disc, which may extend further inward.
For stellar parameters, we adopt $T_{\rm star}=7250$ K, $R_{\rm star}=65$ R, $M_{1}=0.75$ M, and $M_{2}=1.65$ M. For the disc, we adopt an outer radius of $r_{\rm out}=175$ AU and the inner radius of three times the binary orbital separation ($a=1.74$ au), giving $r_{\rm in}=5.22$ au Kluska_2018.
[…]
Corporaal_2023 modelled the surface density using a piecewise power law, with a break at $r_{\rm mid}=3~R_{\rm in}$, and power-law indices of $n_{\rm in}=1.5$ and $n_{\rm out}=-1$ in the inner and outer regions, respectively:
$\Sigma(r)=\Sigma_{0}\left(\frac{r_{\rm mid}}{r}\right)^{n_{\rm in}}\,\text{for}\,r_{\rm in}<r<r_{\rm mid},$ (15)
$\Sigma(r)=\Sigma_{0}\left(\frac{r_{\rm mid}}{r}\right)^{n_{\rm out}}\,\text{for}\,r_{\rm out}>r>r_{\rm mid}.$
The surface density normalisation, $\Sigma_{0}$, is then obtained from the total disc mass using (not valid for $n=2$):
$M_{\rm disc}=\Sigma_{0}\left(\frac{2\pi r_{\rm mid}^{n_{\rm in}}(r_{\rm mid}^{2-n_{\rm in}}-r_{\rm in}^{2-n_{\rm in}})}{(2-n_{\rm in})}+\frac{2\pi r_{\rm mid}^{n_{\rm out}}(r_{\rm out}^{2-n_{\rm out}}-r_{\rm mid}^{2-n_{\rm out}})}{(2-n_{\rm out})}\right)$ (16)
Figure 4 shows the resulting Toomre profiles for both disc mass estimates. If the higher mass inferred by Corporaal_2023 is accurate, the disc reaches $Q<4$ in its outer regions—approaching but not exceeding the fragmentation threshold. This places IRAS 08544-4431 among the most gravitationally unstable post-AGB discs in our sample, though it remains above the threshold for fragmentation. By contrast, the lower mass estimate from Bujarrabal_2018 yields a Toomre profile that stays well within the stable regime across all radii. These results highlight the sensitivity of the Toomre parameter to disc mass assumptions and underscore the need for more robust constraints on gas and dust masses in evolved systems.
Figure 4: The Toomre parameter $Q$ as a function of radius for the post-AGB system IRAS 08544-4431. The shaded region indicates the uncertainty in the total disc mass, based on estimates from Bujarrabal_2018 and Corporaal_2023.
4.2 An Exploration of the Parameter Space for Gravitational Instability Across Different disc Masses and Sizes
Figure 5: Contours showing variation of the Toomre parameter $Q$ as a function of disc radius and mass, evaluated at the outermost radius $r_{\rm out}$ for post-AGB discs with surface density exponent $n=0.5$ and $n=1$ (left and centre panels), and at the innermost radius $r_{\rm in}$ for $n=3$ (right panel). All calculations adopt $R_{\rm star}=100$ R, $T{\rm star}=5000$ K, and $M_{\rm cnt}=1.5$ M.
To complete our investigation, we explore the gravitational stability of discs across a range of mass and size configurations. Specifically, we plot the Toomre parameter $Q$ as a function of disc mass and outer radius to assess whether a given system - such as a post-AGB binary with a more massive disc in the past - could have supported planet formation via gravitational instability.
We consider three central star cases: post-AGB binaries (Figure 5), main-sequence stars (Figure 6, top panel), and white dwarfs (Figure 6, bottom panel). These plots provide a reference framework for gauging the gravitational stability of discs around stars at different evolutionary stages.
In all cases, we adopt a power-law surface density profile (Equation 10) with $n=1$, and additionally show $n=0.5$ for post-AGB systems. For steep profiles ($n=3$), where $Q$ decreases with radius, we evaluate $Q$ at the inner radius ($r_{\rm in}$) instead of $r_{\rm out}$, as shown in the right panel of Figure 5. We adopt fixed parameters of $r_{\rm in}=3$ AU (corresponding to the dynamical truncation radius for a $T_{\rm subl}=1200$ K). The contours represent the value of $Q$ at the disc’s outer edge ($r=r_{\rm out}$), which corresponds to the minimum $Q$ for surface density slopes $n<2$. For $n>2$, where $Q$ decreases outward, the minimum occurs at the inner edge, and we report that value instead.
We find that only sufficiently massive discs are prone to enter the Toomre-unstable regime. Discs with representative post-AGB masses ($\sim 0.01$ M) are unlikely to becomes unstable under present-day conditions and support planet formation via gravitational instability. This is in contrast with discs around main-sequence stars and white dwarfs (Figure 6) which can become gravitationally unstable with much lower disc masses.
We next examine whether the likely conditions of post-AGB discs at the time of their formation would have made them more susceptible to instability and whether planet formation could have occurred during that earlier phase.
[…]
5 Were Discs Less Stable Earlier in Their Lifetimes?
More massive and radially compact discs are more prone to gravitational instability. If the discs we observe today were more massive and compact in the past, they may have been unstable at earlier epochs – even if they appear stable today. Here we calculate the viscous evolution of disks to determine their properties in the past and to determine whether, at that time, they may have been gravitationally unstable.
Viscous spreading naturally causes discs to expand and redistribute mass over time 1974lyndenbell; 1981pringle; 1998hartmann. The evolution of the disc’s surface density under viscous processes is governed by the diffusion equation:
$\frac{\partial\Sigma}{\partial t}=\frac{3}{r}\frac{\partial}{\partial r}\left[r^{1/2}\frac{\partial}{\partial r}\left(r^{1/2}\nu\Sigma\right)\right]$ (17)
where $\nu$ is the kinematic viscosity. Assuming $\nu$ varies with radius as a power law, $\nu\propto r^{\gamma}$ (index is $\gamma$), the radial evolution of the surface density can be solved analytically 1998hartmann:
$\Sigma(r,T)=\frac{C}{3\pi\nu_{0}}\left(\frac{r}{r_{\rm ch,0}}\right)^{-\gamma}T^{-\frac{2.5-\gamma}{2-\gamma}}\exp\left(-\left(\frac{r}{r_{\rm ch,0}}\right)^{2-\gamma}\frac{1}{T}\right)$ (18)
where $C$ is a scaling constant, $\nu_{0}=\nu(r=r_{\rm ch,0})$, $r_{\rm ch,0}$ is the characteristic radius at $t=0$, and $T=1+t/t_{\rm visc}$ is a dimensionless time. The viscous timescale of the disc, $t_{\rm visc}$, is given by:
$t_{\rm visc}=\frac{r_{\rm ch,0}^{2}}{3(2-\gamma)^{2}\nu_{0}}.$ (19)
Equation 18 is the time-dependent analogue of Equation 13 from Section 2.2.2, showing that surface density profiles spread out from initially more centrally concentrated configurations. The characteristic radius evolves as 2009Andrews
$r_{\rm ch}(T)=r_{\rm ch,0}T^{\frac{1}{2-\gamma}}$ (20)
Under the same assumptions, the total disc mass decreases over time due to accretion. 2009Andrews give the mass evolution as:
$M_{\rm disc}(T)=M_{\rm disc,0}T^{\frac{-1}{2(2-\gamma)}},$ (21)
where $M_{\rm disc,0}$ is the initial disc mass. In principle, Equations 20 and 21 can be used to infer past disc properties if the age is known. However, disc ages are poorly constrained for both YSOs and post-AGB systems. Stellar ages cannot be used as reliable proxies for disc ages, particularly in YSOs 2009Andrews.
Here, we assume that post-AGB disc evolution is also dominated by viscous processes, as in YSO discs, although the influence of binary companions may modify this picture. The gravitational torques from companions could affect the disc’s evolution, making these equations less valid 1998hartmann. However, since fragmentation is typically expected in outer disc regions, we assume binary effects are second-order for this analysis.
It is unclear how long discs - even in the case of YSOs - remain in the viscous evolution regime, i.e., how far back in time the analytical relations of Equations 20 and 21 apply. However, the uncertainty in the viscosity parameter and disc age can be mitigated by expressing the evolution of the disc’s mass and radius in terms of the number of viscous timescales that have elapsed, where each viscous timescale depends on the initial radius via Equation 19.
In Figure 7, we plot the Toomre parameter $Q$ as a function of radius for several YSO discs and a representative post-AGB disc ($M_{\rm disc,now}=0.02$ M, $r_{\rm ch,now}=100$ au, and assuming $\gamma=1$), evolved backwards in time for different numbers of elapsed viscous timescales ($T=t/t_{\rm visc}+1$). We assume constant inner radii during viscous evolution. For post-AGB disks, the inner radius is set by the dynamical truncation radius, while for YSOs, it is set by the sublimation radius (taken here as the radius where $T_{\rm subl}=1200$ K). These boundaries depend on stellar properties rather than viscous evolution. Figure 7 shows that even at earlier epochs, some of the discs were likely gravitationally stable —particularly in the cases of PDS70, HD169142, and the post-AGB example. However, in the case of GQ Lup, HL Tau, and Elias 2-27, they could have been gravitationally unstable in their pasts.
Our analysis includes all YSO systems discussed in Section 3, except AB Aurigae and L1448 IRS3B, whose density profiles are not compatible with the form in Equation 13. For post-AGB discs, we assume they follow the viscously evolving form of Equation 13, thus assigning them a characteristic radius. This differs from the simpler power-law model used in earlier sections.
To convert viscous timescales into ages, we adopt the Shakura–Sunyaev Shakura1973 $\alpha$-disc prescription for viscosity ($\nu$):
$\nu=\alpha c_{s}H$ (22)
where $\alpha$ is the viscosity efficiency parameter, and $c_{s}$ and $H$ are the sound speed and scale height, respectively. The scale height $H$ of the disc can be found with the following equation:
$H(r)=\frac{c_{\rm s}(r)}{\Omega(r)}.$ (23)
Typical values of $\alpha$ in planetary discs range from 0.0005 to 0.08, based on observed accretion rates 2009Andrews. Combining this with the expressions for $c_{s}$ and $H$ (Equations 2 and 23) and setting $\eta=1$, the radial dependence of viscosity becomes:
$\nu(r)=\alpha\left(\frac{k_{\rm B}T_{\rm sub}}{2m_{\rm p}}\right)\left(\frac{R_{\rm sub}^{0.75}}{(GM_{\rm cnt})^{0.5}}\right)r^{0.75}$ (24)
Outside the sublimation radius, we evaluate $\nu$ at $r=r_{\rm ch,0}$ to determine $\nu_{0}$. Here, we assume that the stellar parameters have not evolved significantly over the relevant timescales, so the sublimation radius remains approximately constant. For convenience, we group all constants into a single term $K$, allowing us to write:
$\nu_{0}=Kr_{\rm ch,0}^{0.75},$ (25)
yielding a viscous timescale of (for $\gamma=1$):
$t_{\rm visc}=\frac{r_{\rm ch,0}^{1.25}}{3K}$ (26)
Thus, the disc age can be expressed as $t_{\rm visc}\times(T-1)$.
To calculate these parameters, we adopt representative post-AGB values similar to that used in Section 4 and Figure 3. Disc ages for various assumed viscosity parameters ($\alpha$) are presented in Table 2. Importantly, varying $T=t/t_{\rm visc}+1$ alters both the inferred disc age and the initial characteristic radius $r_{\rm ch,0}$, but we find that the total system age remains relatively stable for each specific value of $\alpha$.
We find that the disc age does not vary significantly with changes in the scaled time $T$ for a given $\alpha$, since the viscous timescale itself adjusts accordingly. However, for fixed $T$, the disc age scales linearly with $\alpha$ (for $\gamma=1$). The stellar ages compiled from the literature are included for comparison and should be considered upper limits to the corresponding disc ages. In general, lower $\alpha$ values yield disc ages that exceed the stellar ages—consistent with the findings of 1998hartmann, who argued that $\alpha\approx 0.01$ is typical for YSO discs. For Elias 2-27 specifically, 2009Andrews derived $\alpha=0.03$ from accretion rate measurements, which supports our finding that an assumed $\alpha=0.01$ overestimates the disc age relative to the stellar age.
For post-AGB systems, the inferred ages are consistent with their expected lifetimes of $\sim$$10^{5}$ yr Bujarrabal_2016.
Figure 7: Toomre parameter $Q$ as a function of orbital radius for our YSO sample (excluding AB Aurigae and L1448 IRS3B, for which a characteristic radius is not available) and a representative post-AGB disc (central object parameters similar to those used in Figure 3; a typical value of $\gamma=1$ was assumed due to lack of empirical data). We show $Q$ evaluated at the initial state of each disc for the different assumed viscous timescales.
Table 2: Disc age parameters estimated for each system by assuming different values of the viscosity parameter $\alpha$. For each case, we list the corresponding viscous timescale ($t_{\rm visc}$), number of viscous timescales elapsed ($t/t_{\rm visc}$), inferred disc age ($t$), and the initial characteristic radius ($r_{\rm ch,0}$). For post-AGB discs, we assume $\gamma=1$ due to a lack of empirical values; for YSOs, $\gamma$ is taken from Table 1. See Section  5 for more details.
[…]
- $\alpha$ $_{\rm visc} t/t$$_{\rm visc} tt$ $r_{\rm ch,0}$
- - (yr) - (yr) (au)
typical post-AGB $10^{-2}$ $5.4\times 10^{4}$ 11 $5.4\times 10^{4}$ $50$
$10^{-2}$ $1.4\times 10^{4}$ 55 $6.9\times 10^{4}$ $17$
$10^{-2}$ $6.5\times 10^{3}$ $10$ $6.5\times 10^{4}$ $9.1$
$10^{-3}$ $5.4\times 10^{5}$ 11 $5.4\times 10^{5}$ $50$
$10^{-4}$ $5.4\times 10^{6}$ 11 $5.4\times 10^{6}$ $50$
HL Tauri $10^{-2}$ $7.3\times 10^{5}$ 11 $7.3\times 10^{5}$ $58$
$10^{-2}$ $3.9\times 10^{5}$ 55 $2.0\times 10^{6}$ $35$
$10^{-2}$ $2.8\times 10^{5}$ $10$ $2.8\times 10^{6}$ $27$
$10^{-3}$ $7.3\times 10^{6}$ 11 $7.3\times 10^{6}$ $58$
$10^{-4}$ $7.3\times 10^{7}$ 11 $7.3\times 10^{7}$ $58$
Elias 2-27 $10^{-2}$ $8\times 10^{6} 11$ $1.8\times 10^{6}$ $100$
$10^{-3}$ $8\times 10^{7} 11$ $1.8\times 10^{7}$ $100$
$10^{-4}$ $8\times 10^{8} 11$ $1.8\times 10^{8}$ $100$
PDS 70 $10^{-2}$ $4.9\times 10^{5}$ 11 $4.9\times 10^{5}$ 2020
$10^{-3}$ $4.9\times 10^{6}$ 11 $4.9\times 10^{6}$ 2020
$10^{-4}$ $4.9\times 10^{7}$ 11 $4.9\times 10^{7}$ 2020
HD 169142 $10^{-2}$ $6.9\times 10^{5}$ 11 $6.9\times 10^{5}$ $50$
$10^{-3}$ $6.9\times 10^{6}$ 11 $6.9\times 10^{6}$ $50$
$10^{-4}$ $6.9\times 10^{7}$ 11 $6.9\times 10^{7}$ $50$
GQ Lupi $10^{-2}$ $8\times 10^{6} 11$ $1.8\times 10^{6}$ $68$
$10^{-3}$ $8\times 10^{7} 11$ $1.8\times 10^{7}$ $68$
$10^{-4}$ $8\times 10^{8} 11$ $1.8\times 10^{8}$ $68$
  • The estimated ages for the central stars of these discs are 1 Myrs for HL Tauri Booth_2020, 1 Myrs for Elias 2-27 2009Andrews, 5.4 Myrs for PDS 70 Keppler_2019, 6 Myrs for HD 169142Hammond_2023, and 2-5 Myrs for GQ Lupi Wu_2017.
5.1 The Effects of Varying the Temperature Profile
In the analysis so far, we have assumed that the disc temperature follows the radial dependence given by Equation 7. However, the temperature gradient is linked to the viscous disc’s surface density index $\gamma$ in Equation 18. Specifically, if $T(r)\propto r^{p}$, and $\nu\propto r^{\gamma}$, then $\gamma=1.5+p$ 1998hartmann.
In reality, the assumption of constant $\alpha$ may not hold, and the Shakura-Sunyaev prescription may not strictly apply. Therefore, we have not varied the temperature profile as a function of the observationally derived $\gamma$ values in our main analysis. However, in this section we explore how changes in the temperature index $p$ affect the gravitational stability of a post-AGB disc, using the corresponding value of $\gamma=1.5+p$ in each case.
[…]
Figure 8 shows the resulting $Q$ profiles. As expected, steeper temperature profiles (larger $p$) lower the $Q$ parameter across the disc. However, even for the steepest case considered, the disc remains gravitationally stable. We emphasise that this variation does not include different temperature slopes across the sublimation radius, as applied in earlier figures.
Figure 8: The Toomre parameter $Q$ as a function of orbital radius for a representative post-AGB disc for varying temperature profile index, $p$, and the corresponding $\gamma$, assuming $\gamma=1.5+p$. The central binary properties are similar to Figure 3, while assuming the surface density profile of a viscous disc (Equation 13). We adopt $r_{\rm ch}=100$ au and an inner dynamic truncation radius of $r_{\rm in}=3$  au.
6 Planet Formation in Post-Common Envelope Discs
In this section we apply our gravitational instability framework to the post-common envelope binary NN Serpentis, where the presence of two proposed circumbinary planets has been interpreted as evidence for second generation planet formation in a remnant disc. These planets were first inferred from eclipse timing variations Beuermann_2010. Given their proximity to the central binary, it has been argued that they could not have formed during the system’s main sequence evolution and subsequently survived both the giant phase and the common-envelope expansion. Instead, it has been proposed that they formed within acircumbinary disc created during the common-envelope interaction.
The existence of these circumbinary planets, however, has been the subject of debate. While the planets were first inferred from eclipse timing variationsBeuermann_2010; Beuermann_2013, and later studies assessed the dynamical stability of the proposed orbits, suggesting that such configurations could be stable 2012Horner; Marsh_2013, their interpretation as planets is regarded with some caution. In particular, Marsh_2013 noted that the planets are only indirectly detected, and that alternative mechanisms – most notably magnetic activity cycles (Applegate’s mechanism) – can also reproduce the observed eclipse timing variations.
On the premise of the planets’ existence, 2014schleicher developed an analytical framework to estimate the properties of a circumbinary disk formed through common-envelope ejection. By applying gravitational instability theory, they demonstrated that the planetary masses and orbital separations inferred for NN Ser are consistent with formation in such an environment.
We apply our updated gravitational instability model to determine whether such a disk is indeed prone to making planets. We carry out our analysis for a disk mass like the one used by 2014schleicher, and for a disk mass that is 10 times smaller. The reason for this is that while some circumbinary gas has been detected around some post-CE binaries Li2025, recent theoretical developments have suggested that full envelope ejection is likely and less massive disks would then form by fall-back of a small amount of bound material 2015ivanova; 2020Reichardt.
In Figure 9 we show the radial variation of the Toomre parameter $Q$ for the two assumed disk masses, 0.1 M and 0.01 M. The central source is a white dwarf with with $M_{c}=0.535$ M, $R_{\rm star}=0.0189$ R, and $T_{\rm star}=57\,000$ K 2004Haefner. We have also assumed an outer circumbinary disk radius of $r_{\rm out}=100$ au (similar to Figure 3, with the only difference being that the central object is a white dwarf instead of a post-AGB star). As expected, more massive discs are more prone to gravitational instability. Notably, despite its high temperature, the compact size of the white dwarf results in a lower overall luminosity and, consequently, a cooler mid-plane disk temperature, compared to systems with larger, even if cooler, post-AGB stars. This reduced thermal support lowers the sound speed, thereby decreasing $Q$ and making the disc more susceptible to fragmentation. However, even with these more favourable conditions, only the more massive disk leads to possible planet formation by gravitational instability.
While 2014schleicher assumed that the entire disc around NN Ser is marginally unstable ($Q(r)\sim 1$), we now investigate whether this assumption holds when a more realistic temperature profile than the one they assumed is adopted. To do so, we use Equation 7, incorporating the observed properties of the central white dwarf as listed above. Following the framework proposed by 2014schleicher, we find an outer radius of $r_{\rm out}=7.50$ au for the disc, which is slightly different from the value they found ($r_{\rm out}=6.48$ au – this difference is further discussed in B). We also use their values of $M_{\rm disk}=0.146$M, a power-law index of $n=1$ for the surface density profile (Equation 10), and a mass of the central object of $M_{\rm core}+M_{2}=0.646$ M (see Table 3).
Figure 10 shows the radial variation of the Toomre parameter, $Q$, for NN Ser using their proposed disk mass, a disk radius of $r_{\rm out}=7.50$ au and the temperature profile of Equation 7. The figure shows that even with a more realistic thermal structure, the disc becomes gravitationally unstable beyond $\sim 3$ au - precisely where the proposed planets are located. This analysis therefore confirms that the gravitational instability mechanism remains viable under more physically motivated conditions than those selected by 2014schleicher, but only if the disk mass is on the high, rather than low side (0.1 Mvs. 0.01 M (for further details see B).
Table 3: Parameters adopted or determined by 2014schleicher in their study of NN Ser. For details see B.
[…]
Disk Mass, $M_{\rm disk}$ 0.146 M
Ejected Mass, $M_{\rm ej}$ 1.24 M
Primary mass, $M_{1}$ 2 M
Companion mass, $M_{2}$ 0.111 M
Primary core mass, $M_{\rm c}$ 0.535 M
Primary radius, $R_{*}$ 185 R
Disc’s surface density index, $n$ 1
Disc’s outer radius, $r_{\rm out}$ 7.5 R [^11] The value we find for $r_{\rm out}$ is different from that of 2014schleicher, see B for more details.
Disc’s ang. mom., $L_{\rm dep}$ $1.2\times 10^{52}$ g cm s-1121212The value we find for $L_{\rm dep}$ is different from that of 2014schleicher, see B for more details.
Disc’s ang. mom. enhancement factor, $\alpha_{\rm L}$ 12.5
critical viscosity, $\alpha_{\rm crit}$ 0.3
Figure 9: The Toomre parameter $Q$ as a function of radius for discs with masses of 0.01 M (left panel) and 0.1 M (right panel) and parameters similar to Figure 3, but with a white dwarf as the central star ($M_{c}=0.535$ M, $R_{\rm star}=0.0189$ R, and $T_{\rm star}=57,000$ K), and an outer disc radius of $r_{\rm out}=100$ au.
Figure 10: The Toomre parameter, $Q$, as a function of orbital radius for the progenitor disc of NN Serpentis, based on the values of 2014schleicher, which are summarised in Table 3. The white dwarf parameters were $T_{\rm eff}=57,000$ K and $R_{\rm star}=0.0189$ R, and the mass of the central binary was set to be $M_{\rm c}+M_{2}=0.646$ M. We note that the outer radius of the disc used to reproduce this plot was $r_{\rm out}=7.50$ au, which we recomputed using the value we find for angular momentum in B.
7 Conclusion
In this study, we have carried out a comprehensive analytical investigation into whether post-AGB circumbinary discs can support planet formation via gravitational instability. Our key findings are:
  1. 1.
    Under realistic assumptions of disc mass 0.020.050.02-0.05M, disc structure, and density and temperature profiles, post-AGB discs are generally gravitationally stable at present. Even systems with the most optimistic parameters—such as IRAS 08544-4431 with a disc mass of 0.10.1M—remain above the fragmentation threshold, with Toomre $Q>1$ throughout the disc, largely due to thermal support from the luminous central star. In contrast, discs around lower-luminosity central objects (e.g. white dwarfs or Sun-like stars) can reach the gravitationally unstable regime at lower masses.
  2. 2.
    Using self-similar viscous disc evolution models, in which the surface density profile retains its shape while the overall scale (mass and characteristic radius) evolves due to viscous spreading, we explored whether post-AGB discs may have been unstable earlier in their lifetimes. We found that, even after evolving them backward in time across several viscous timescales, most remain stable, unlike some protoplanetary discs (e.g. HL Tauri, Elias 2-27, GQ Lupi), which likely passed through an unstable phase during earlier epochs. Our analysis also shows that higher viscosity parameters ($\alpha\sim 10^{-2}$) are more consistent with expected post-AGB disc lifetimes.
  3. 3.
    Revisiting the post-common envelope binary NN Serpentis, we confirm that if the disc parameters proposed by 2014schleicher are assumed, the disc would become gravitationally unstable beyond  $\sim\,3$ au. However, the planet masses predicted by their gravitationally induced clump growth model slightly underpredict those of the observed planet candidates, and recent developments in common envelope theory suggest that fallback discs may be far less massive than previously assumed.
It is important to note that using the Toomre parameter as a diagnostic for gravitational instability, while useful, does not take factors such as vertical structure, turbulence, and binary-driven dynamics into account. Post-AGB disks can differ from YSO disks in these regards as well, hence the results of this study should be treated as a first order analysis.
Overall, our analysis demonstrates that gravitational instability is highly unlikely to account for second-generation planet formation in typical post-AGB discs, as fragmentation would require unrealistically high disc masses ($\gtrsim 0.4$ M), even under compact and cool conditions. The observed signatures of planets or planet-like companions in several post-AGB systems therefore point towards alternative formation channels—most likely a modified version of core accretion operating on shorter timescales in evolved environments. This study provides a first theoretical baseline for interpreting such systems. While future observations will better constrain disc masses, gas-to-dust ratios, temperatures, sizes, and lifetimes—and may uncover additional evidence for planet formation—we will, in parallel, explore additional theoretical models to investigate planet formation pathways in post-main-sequence environments.
Acknowledgments
AP acknowledges the financial support provided by the International Macquarie Research Excellence Scholarship (iMQRES) program, received throughout the duration of this research. DK, ODM and MW acknowledge the support from the Australian Research Council Discovery Project DP240101150.
[…]
Appendix A Parameters of YSO systems
In this section we present the studies we have used to compile the YSO disc properties into Table 1.
We note that the reported inner radii in the literature typically refer to the inner radius of the dust component. Some studies explicitly report an inner gas cavity (such as HD169142), but others do not distinguish between the two. If a study has not stated that the inner radius they have reported is the dust radius or gas radius, we have taken that as the gas inner radius. In the case of some studies that carried out simulations and had reported the inner radius they had taken for the disk cavity, such as AB Aurigae, and Elias 2-27, we used the value they adopted. For PDS 70 there were no inner cavities in literature, hence we adopted $r_{\rm in}=0$.
There are two different reported masses for the disc of AB Aurigae, each obtained with different methods, and we have carried out Toomre analysis with both values. rivier2024 report $3.2\times 10^{-3}~\rm M_{\odot}$ which is based on ALMA and VLA observations to find the dust mass and then use a gas-to-dust ratio to estimate the total disc mass. On the other hand, 2023Speedie have estimated the disc mass based on their observation and subsequent modelling of the gravitational wiggle they observe in the disc, which they argue is a signature of gravitational instability, and they find a disc mass of $0.7~\rm M_{\odot}$ best fits their model. Other than the disc mass, the rest of the tabulated values are those that 2023Speedie have taken for their hydrodynamics simulation (note that this is the only system in this table which uses a power law for surface density rather than Equation 13).
For HL Tauri, the parameters for the star’s radius and luminosity are from 1999menshchikov. The star mass was calculated in 2016pinte by fitting models to CO and HCO+ velocities obtained from ALMA, and the rest of the parameters, most notably the disc mass, were taken from Booth_2020, who have used $^{13}\rm C^{17}\rm O$ measurements from ALMA to model and find the disc mass which is significantly more massive compared to previous measurements.
For Elias 2-27 , we adopted the value of the inner radius of the disc from Paneque_Carre_o_2021, the photospheric temperature and radius of the star from 2009Andrews and the rest from Veronesi_2021, who used ALMA observations to find the disc and star masses. It is noteworthy to mention that this system has been also analyzed to see whether it is prone to instabilities in Perez_2016, albeit with different values, and for their standard opacity value their resulting Toomre estimates are similar to us, noting the slightly different values they have adopted for the disc parameters.
The values for PDS 70 are available in Keppler_2019, from ALMA measurements. We assume the inner radius of the disc to be approximately zero, $r_{\rm in}=0$ au .
For HD 169142, we take the values of star mass, radius, and temperature from Quanz_2013. Quanz_2013 report 3 different temperatures, and the adopted one was calculated from high-resolution spectroscopy by 2006guimaraes who find $T_{\rm eff}=7500$K. We take the characteristic disc radius and other disc parameters from Fedele_2017, who used a best fit to ALMA observations to estimate the disc parameters. For the disc mass, Fedele_2017 mentioned that the total dust mass is $10^{-4}$ M and the best fit they find for gas-to-dust mass ratio is 80, which gives the disc mass to be 0.008 M. They also determined $\Sigma_{\rm ch}=6.5$ g/cm2 as the surface density at the characteristic radius.
For GQ Lup, both disc and star parameters had been summarized by Wu_2017. The disc surface density was fitted using ALMA data. Similarly for L1448 IRS3B, the disc and star parameters have been taken from Reynolds_2021, who used ALMA data.
Appendix B Planet Formation Around Post-Common Envelope Binary Systems
For completeness and consistency we calculate here the mass and orbital parameters of the two planets using a realistic temperature profile that allows us not to assume a value of $Q$, and compare our values with those of 2014schleicher.
2014schleicher started their analysis of the NN Ser post-common envelope (CE) binary and its putative planets, by motivating their choice of disk mass ($\sim 0.1$ M). They then estimated the amount of angular momentum deposited by the companion into the envelope of the AGB star, $L_{\rm dep}$:
$L_{\rm dep}=M_{2}R_{*}\sqrt{\frac{GM_{1}}{R_{*}}}$ (27)
where $M_{1}$ and $R_{*}$ are the AGB star’s mass and radius, respectively and $M_{2}$ is the companion’s mass (see Table 3). In order to account for additional angular momentum that the envelope may already have had at the time of CE in-spiral, they assumed that the actual disk’s specific angular momentum would be a factor of a few larger than the one deposited by the companion. They parameterised this in the following way:
$\frac{L_{\rm disc}}{M_{\rm disc}}=\alpha_{\rm L}\frac{L_{\rm dep}}{M_{\rm ej}}$ (28)
where $\alpha_{\rm L}$ is the enhancement factor for the specific angular momentum, $L_{\rm disc}$ is the angular momentum of the disc and $M_{\rm ej}$ is the mass ejected during the CE phase, but that presumably remains bound and forms the disk.
After the disc has formed, they assumed that its surface density, $\Sigma$, settles into a power-law radial profile with a truncation radius, $r_{\rm out}$, described by Equation 10. If they calculate the angular momentum of the disc assuming a Keplerian velocity surrounding a stripped core with mass $M_{\rm c}$ and a companion star of mass $M_{2}$, they could find the outer radius, $r_{\rm out}$:
$r_{\rm out}=\left(\frac{2.5-n}{2-n}\right)^{2}\left(\frac{L_{\rm disc}}{M_{\rm disc}}\right)^{2}\frac{1}{G(M_{\rm c}+M_{2})},$ (29)
with $L_{\rm disc}/M_{\rm disc}$ from Equation 28.
2014schleicher made the assumption that this disk is gravitationally unstable ($Q=1$). From an initial clump mass, $M_{\rm cl}$, approximately estimated as:
$M_{\rm cl}(r)\approx\Sigma(r)h(r)^{2},$ (30)
the clumps will then grow on a dynamical timescale to a final mass:
$M_{\rm f}(r)=M_{\rm cl}(r)\left(12\pi^{2}\frac{\alpha_{\rm crit}}{0.3}\frac{r}{H(r)}\right)^{0.5},$ (31)
with $\alpha_{\rm crit}$ the critical viscosity parameter. This equation is obtained from the gap-opening mass expression of 1986lin [^13] Equation 34 of 2014schleicher has a missing factor of $\sqrt{\pi}$ compared to that of 1986lin.. They obtain masses of 6.6 MJup and 1.9 MJup (cf. observed values of 7 MJup and 1.7 MJup) for orbital separations of 5.4 au and 3.4 au.
We have summarised the values used by 2014schleicher in Table 3. Using these values, we found $L_{\rm dep}=1.3\times 10^{52}$ erg cm2 s-1 and $r_{\rm out}=7.50$ au (while 2014schleicher reported $1.2\times 10^{52}$ erg cm2 s-1 and $r_{\rm out}=6.48$ au).
We then estimate the final clump mass (Equations 30 and 31) by using a realistic temperature profile and not assuming that $Q=1$ (Equation 7). The predicted final clump masses for orbital separations of 5.4 au and 3.4 au (blue line) are 2 MJup and 1.2 MJup, somewhat lower than the values obtained by 2014schleicher. Despite this small discrepancy, the analysis of 2014schleicher is consistent with what we have found so that we can conclude that if the disk mass is close to their adopted 0.1 M then indeed this disk could make the observed planets by gravitational fragmentation.
Figure 11: Clump masses and final (planet) masses as a function of orbital distance for NN Serpentis using the values used by 2014schleicher, but assuming the temperature profile of Equation 7 and $r_{\rm out}=7.5$ au.